# Robust Cross-View Geo-Localization via Content-Viewpoint Disentanglement

### v1

@@ Abstract
Cross-view geo-localization (CVGL) aims to match images of the same geographic location captured from different perspectives, such as drones and satellites. Despite recent advances, CVGL remains highly challenging due to significant appearance changes and spatial distortions caused by viewpoint variations. Existing methods typically assume that cross-view images can be directly aligned within a shared feature space by maximizing feature similarity through contrastive learning. Nonetheless, this assumption overlooks the inherent conflicts induced by viewpoint discrepancies, resulting in extracted features containing inconsistent information that hinders precise localization. In this study, we take a manifold learning perspective and model the feature space of cross-view images as a composite manifold jointly governed by content and viewpoint information. Building upon this insight, we propose CVD, a new CVGL framework that explicitly disentangles content and viewpoint factors. To promote effective disentanglement, we introduce two constraints: (i) An intra-view independence constraint, which encourages statistical independence between the two factors by minimizing their mutual information. (ii) An inter-view reconstruction constraint that reconstructs each view by cross-combining content and viewpoint from paired images, ensuring factor-specific semantics are preserved. As a plug-and-play module, CVD can be seamlessly integrated into existing geo-localization pipelines. Extensive experiments on four benchmarks, i.e., University-1652, SUES-200, CVUSA, and CVACT, demonstrate that CVD consistently improves both localization accuracy and generalization across multiple baselines.
11footnotetext: Corresponding author: wangdi@xidian.edu.cn
@@ 1 Introduction
The widespread deployment of autonomous vehicles, drones, and other intelligent systems has placed growing demands on navigation and localization technologies. To ensure safe and reliable operation, high-precision positioning services have become a fundamental requirement. Cross-view geo-localization (CVGL), an onboard technique independent of external communication infrastructure, offers a promising solution by estimating absolute geospatial coordinates in the absence of conventional localization signals (e.g., GPS) [18, 3, 25, 8]. Given a query image, typically captured from a ground or drone perspective, the goal is to find a matching satellite image from a georeferenced database to infer the query’s location. Most existing approaches formulate CVGL as an image retrieval task [17, 23], training deep neural networks (DNNs) to learn visual similarity across different views. However, the viewpoint disparity between query and satellite imagery introduces severe spatial distortions and appearance variations, making robust matching inherently challenging.
@@
Figure 1: Comparison between previous methods and our CVD. Left: Existing methods can be interpreted as operating on a single manifold $\mathcal{M}$, where contrastive objectives directly pull positive pairs closer and push negative pairs away. Right: Our method learns disentangled representations by mapping inputs onto two submanifolds corresponding to content $\mathcal{M}_{c}$ and viewpoint $\mathcal{M}_{v}$. This separation is enforced via two constraints (see Section 3.3 and Section 3.4), promoting effective disentanglement and thereby enhancing cross-view matching performance.
@@
Recent efforts aim to alleviate the viewpoint discrepancy between query and satellite images. A common strategy is to apply predefined geometric transformations to satellite images, such as polar transformation or orthorectification, aligning their spatial layout with that of the query view. However, the effectiveness of these methods relies on prior knowledge of the geometric relationship between views, and may degrade when the query is not spatially centered within the satellite image [31].
@@
Despite recent advances in CVGL, many existing methods [9, 22, 31, 4, 27] still follow a common training paradigm, i.e., directly applying contrastive learning to pull the features of positive pairs closer and push those of negative pairs away. While various architectural or training refinements have been proposed, these methods largely overlook the semantic inconsistencies introduced by drastic viewpoint differences. These inconsistencies disrupt the alignment between positive pairs and ultimately limit the cross-view matching performance.
@@
In this paper, we revisit the CVGL task from a manifold learning perspective by modeling the feature space of cross-view images as a composite manifold jointly governed by content and viewpoint factors. As illustrated in Figure 1 (left), prior methods can be interpreted as learning representations on a composite manifold $\mathcal{M}$, where both factors are inherently entangled. Such entanglement introduces viewpoint-induced conflicts into the learned representations, thereby undermining the robustness of contrastive alignment. To address this limitation, we propose CVD (Content-Viewpoint Disentanglement), a generic CVGL framework that explicitly factorizes the feature space into two submanifolds: a content submanifold and a viewpoint submanifold (see Figure 1 (right)). The content encodes view-agnostic structural information, while the viewpoint captures view-specific appearance variations. As shown in Figure 2 (Training Stage), CVD adopts an embed-disentangle-reconstruct paradigm. Each image is first embedded into a shared feature space, then projected onto independent content and viewpoint submanifolds, and finally recombined in an image reconstruction task. To facilitate effective disentanglement, we impose two dedicated constraints: an intra-view independence constraint and an inter-view reconstruction constraint. The former encourages statistical independence between content and viewpoint by minimizing their mutual information, while the latter ensures that each factor retains its intended semantics by reconstructing images using the content of one view and the viewpoint of the other. In addition, we apply a standard contrastive loss (e.g., InfoNCE) to align content representations of matched pairs across views.
To the best of our knowledge, CVD is the first CVGL framework to explicitly disentangle content and viewpoint information. Unlike existing methods that rely on architectural redesign, CVD serves as a plug-and-play module that can be seamlessly integrated into existing pipelines. It introduces no additional cost during inference while substantially improving matching accuracy and robustness. In summary, our contributions are as follows:
@@
  • We revisit the cross-view geo-localization (CVGL) task from a manifold learning perspective and propose CVD, the first framework that explicitly disentangles content and viewpoint to suppress viewpoint-induced conflicting information.
@@
  • We introduce two constraints to facilitate effective disentanglement: an intra-view independence constraint that facilitates the independence between content and viewpoint, and an inter-view reconstruction constraint that ensures each factor retains its intended semantics.
@@
  • CVD can be seamlessly integrated into existing pipelines, incurs no additional inference overhead, and consistently improves cross-view matching performance and generalization across multiple challenging benchmarks.
@@ 2 Related Work
CVGL has witnessed significant progress with the availability of large-scale geo-tagged datasets and advances in deep learning. Existing methods predominantly follow a Siamese-based framework and can be broadly categorized into three research directions: data augmentation strategies, architectural innovations, and feature representation learning.
2.1 Data Augmentation Strategies
@@
Data augmentation has become a widely adopted strategy in CVGL tasks [19, 12, 26, 6]. To address cross-view misalignment, Liu et al.[9] incorporate camera orientation as auxiliary input, while SAFA[15] applies a polar transformation to align aerial and ground-level panoramas. Several methods further improve robustness by diversifying input image appearances. For instance, Lv et al.[10] simulate extreme weather conditions by injecting synthetic variations into University-1652. More recently, training-aware data sampling has emerged as a complementary strategy. Sample4Geo[4] introduces two curriculum-driven strategies: one leveraging geographically adjacent samples for easier early-stage alignment, and another mining hard negatives to refine the decision boundary. Game4Loc [5] proposes a mutual-exclusion sampling mechanism that enforces strict decorrelation between positive and negative pairs, thereby enhancing contrastive supervision in cross-view matching.
2.2 Backbone Innovations
@@
In parallel, a line of work focuses on designing more powerful visual backbones to enhance localization performance [13, 29]. For example, L2LTR [22] exploits self-attention to model long-range dependencies, effectively reducing visual ambiguity in cross-view. RK-Net [7] introduces a lightweight unit-difference attention module that enables joint learning of dense features and salient keypoints, without requiring additional annotations. SAIG [33] proposes an efficient backbone tailored for CVGL by replacing the MLP blocks in standard Transformers with spatially-aware mixing layers and low-dimensional projections, yielding a more compact and structured representation.
2.3 Feature Representation Learning
@@
Many studies focus on learning more effective visual representations to enhance cross-view matching [32, 11, 30, 24]. A common strategy involves refining alignment mechanisms between views. For instance, Shi et al.[16] proposed a dynamic similarity matching network to estimate directional alignment, thereby reducing cross-view discrepancies. FSRA[2] leverages transformer-based heatmaps to perform region-level alignment, while LPN [20] incorporates contextual cues via a square-ring partitioning strategy to improve part-based representations. SDPL [1] builds upon LPN by introducing a shifting-fusion mechanism to generate multiple complementary part sets, which are then adaptively aggregated to enhance robustness against spatial shifts and scale variations. Several methods also incorporate inductive priors modeling to enhance feature expressiveness. TransGeo [31] adopts a non-uniform cropping strategy that discards low-information regions while reallocating resolution to semantically salient areas, enhancing accuracy without increasing computational cost. GeoDTR [27] incorporates spatial layout priors through a geometry-aware extraction module, combined with counterfactual training to exploit structural information. MCCG [14] enriches feature diversity by jointly modeling spatial and channel-wise attentions.
@@
Although prior methods perform well across various scenarios, many rely on auxiliary components that increase inference-time overhead. In contrast, we propose a plug-and-play training paradigm that explicitly disentangles content and viewpoint from raw features, enhancing cross-view correspondence without introducing additional parameters or latency at inference.
@@
3 Methodology
3.1 Problem Formulation
Considering a set of image pairs $\{(\mathcal{I}_{i}^{\mathrm{q}},\mathcal{I}_{i}^{\mathrm{s}})\}_{i=1}^{N}$, where superscripts $\mathrm{q}$ and $\mathrm{s}$ denote query (e.g., drone or ground) and satellite images, respectively, and $N$ is the number of pairs. Each pair depicts the same geographic location. In the CVGL task, given a query image $\mathcal{I}_{q}^{\mathrm{q}}$ with index $q$, the objective is to retrieve its best-matching satellite image $\mathcal{I}_{b}^{\mathrm{s}}$ from the georeferenced database, where $b\in\{1,\ldots,N\}$.
For effective comparison between cross-view images, we aim to modulate raw representations by explicitly disentangling two factors: content and viewpoint. Concretely, we model the feature space $\mathcal{Z}$ as a representation of composite manifold $\mathcal{M}$ structured by two independent submanifolds, $\mathcal{M}_{c}$ and $\mathcal{M}_{v}$, corresponding to content and viewpoint, respectively. Each feature representation is viewed as a sample from two latent random variables, $C\sim p(C)$ and $V\sim p(V)$, defined over $\mathcal{M}_{c}$ and $\mathcal{M}_{v}$. These are composed via a function $f:\mathcal{M}_{c}\times\mathcal{M}_{v}\rightarrow\mathcal{Z}$ that maps both factors to a point in the feature space. Assuming statistical independence between the factors, i.e., $p(C,V)=p(C)\,p(V)$, the resulting distribution over $\mathcal{Z}$ is given by the push-forward measure $f_{\#}(p(C)\times p(V))$. This assumption is well-aligned with the CVGL task, where identical scenes may be observed under diverse perspectives.
To realize the above formulation, we design manifold encoders that decompose each image $\mathcal{I}^{\mathrm{u}}$ into two distinct representations: a content embedding $f_{c}(\mathcal{I}^{\mathrm{u}})\in\mathcal{M}_{c}$ and a viewpoint embedding $f_{v}(\mathcal{I}^{\mathrm{u}})\in\mathcal{M}_{v}$. Cross-view matching is subsequently performed in the $\mathcal{M}_{c}$ by retrieving the nearest neighbor of a query image $\mathcal{I}_{q}^{\mathrm{q}}$ via its content embedding:
$b=\underset{i\in\{1,\ldots,N\}}{\arg\min}d(f_{c}(\mathcal{I}_{q}^{\mathrm{q}}),f_{c}(\mathcal{I}_{i}^{\mathrm{s}})),$ (1)
where $d(\cdot,\cdot)$ denotes a distance metric. For notation compactness, we will use superscript $u$  [^2] We adopt this convention throughout the paper. for cases that apply to both query ($\mathrm{q}$) and satellite ($\mathrm{s}$) views.
3.2 Proposed Methodology
As illustrated in Figure 2, CVD adopts a Siamese architecture consisting of two symmetric branches for the query and satellite views. Each branch comprises three sequential components: manifold embedding, information disentanglement, and cross-reconstruction.
Manifold Embedding. Each input image $\mathcal{I}^{\mathrm{u}}$ ($u\in\{\mathrm{q},\mathrm{s}\}$) is first processed by manifold encoders $E$, yielding a raw $d$-dimensional feature representations $\mathbf{z}^{\mathrm{u}}=E(\mathcal{I}^{\mathrm{u}})\in\mathbb{R}^{d}$. Owing to the nature of DNN encoders, the distribution of $\mathbf{z}^{\mathrm{u}}$ can be viewed as residing on a composite manifold jointly governed by content and viewpoint information. Since our primary focus is on the training paradigm, we adopt the same DNN encoders as those used in the respective baselines to ensure fair and consistent comparisons.
@@ 3.2 Proposed Methodology
Information Disentanglement. Once we obtain the feature representation $\mathbf{z}^{\mathrm{u}}$, we project it into two statistically independent components using two parallel $3\times 3$ convolutional layers with a channel ratio of $\alpha$. One is used to represent the content embedding, denoted as $\mathbf{z}_{c}^{\mathrm{u}}=f_{c}(\mathbf{z}^{\mathrm{u}})\in\mathbb{R}^{\alpha d}$, which provides view-agnostic scene structure. The other is employed for the viewpoint embedding, denoted as $\mathbf{z}_{v}^{\mathrm{u}}=f_{v}(\mathbf{z}^{\mathrm{u}})\in\mathbb{R}^{(1-\alpha)d}$, capturing view-specific attributes. To promote effective disentanglement, we introduce an intra-view independence constraint (Section 3.3) that minimizes the mutual information $\mathrm{MI}(Z_{c}^{\mathrm{u}};Z_{v}^{\mathrm{u}})$, where $Z^{\mathrm{u}}$ denotes the random variables of $\mathbf{z}^{\mathrm{u}}$, thereby encouraging statistical independence between content and viewpoint factors.
@@
Cross-Reconstruction. While the independence constraint enhances factor separation, it may inadvertently lead to degenerate solutions or information loss. To mitigate this, we introduce an inter-view reconstruction constraint that encourages each factor to retain its intended information through cross-view image reconstruction. Specifically, we train two decoders, $D^{\mathrm{q}}$ and $D^{\mathrm{s}}$, to perform bidirectional reconstruction between paired views by swapping content and viewpoint embeddings, i.e., reconstructing each image using its own viewpoint and the content of the other. By enforcing accurate reconstruction from these hybrid embeddings, the model is incentivized to encode factor-specific information in each representation. This cross-view supervision not only prevents information collapse but also reinforces disentanglement. In the following, we describe the two specific constraints in CVD.

### v2

# Robust Drone-View Geo-Localization via Content-Viewpoint Disentanglement

@@ Abstract
Drone-view geo-localization (DVGL) aims to match images of the same geographic location captured from drone and satellite perspectives. Despite recent advances, DVGL remains challenging due to significant appearance changes and spatial distortions caused by viewpoint variations. Existing methods typically assume that drone and satellite images can be directly aligned in a shared feature space via contrastive learning. Nonetheless, this assumption overlooks the inherent conflicts induced by viewpoint discrepancies, resulting in extracted features containing inconsistent information that hinders precise localization. In this study, we take a manifold learning perspective and model the feature space of cross-view images as a composite manifold jointly governed by content and viewpoint. Building upon this insight, we propose CVD, a new DVGL framework that explicitly disentangles content and viewpoint factors. To promote effective disentanglement, we introduce two constraints: (i) an intra-view independence constraint that encourages statistical independence between the two factors by minimizing their mutual information; and (ii) an inter-view reconstruction constraint that reconstructs each view by cross-combining content and viewpoint from paired images, ensuring factor-specific semantics are preserved. As a plug-and-play module, CVD integrates seamlessly into existing DVGL pipelines and reduces inference latency. Extensive experiments on University-1652 and SUES-200 show that CVD exhibits strong robustness and generalization across various scenarios, viewpoints and altitudes, with further evaluations on CVUSA and CVACT confirming consistent improvements.
@@ 1 Introduction
Figure 1: Comparison between previous methods and our CVD. Left: Existing methods can be interpreted as operating on a single manifold $\mathcal{M}$, where contrastive objectives directly pull positive pairs closer and push negative pairs away. Right: Our method learns disentangled representations by mapping inputs onto two submanifolds corresponding to content $\mathcal{M}_{c}$ and viewpoint $\mathcal{M}_{v}$. This separation is enforced via two constraints (see Sec. 3.3 and Sec. 3.4), promoting effective disentanglement and thereby enhancing cross-view matching performance.
The widespread deployment of drones and other intelligent systems has placed growing demands on navigation and localization technologies. To ensure safe and reliable operation, high-precision positioning services have become a fundamental requirement. Drone-view geo-localization (DVGL), an onboard technique independent of external communication infrastructure, offers a promising solution by estimating absolute geospatial coordinates in the absence of conventional localization signals (e.g., GPS) [sun2023cross, dai2023vision, zhang2024aligning, lin2015learning]. Given a drone image, the goal is to find a matching satellite image from a georeferenced database to infer the drone’s location. Most existing approaches formulate DVGL as an image retrieval task [sun2021multisensor, ye2024cross2], training deep neural networks (DNNs) to learn visual similarity across different views. However, the viewpoint disparity between drone and satellite imagery introduces severe spatial distortions and appearance variations, making robust matching inherently challenging.
@@
Recent efforts aim to alleviate the viewpoint discrepancy between drone and satellite images. A common strategy is to employ predefined geometric transformations to satellite images, such as polar transformation or orthorectification, aligning their spatial layout with that of the drone view. However, the effectiveness of these methods relies on prior knowledge of the geometric relationship between the two views and may degrade when the drone image is not spatially centered within the satellite image [zhu2022transgeo].
@@
Despite recent advances in DVGL, many existing methods [liu2019lending, yang2021cross, zhu2022transgeo, deuser2023sample4geo, zhang2023cross] still follow a common training paradigm, i.e., directly applying contrastive learning to pull the features of positive pairs closer and push those of negative pairs away. While various architectural or training refinements have been proposed, these methods largely overlook the semantic inconsistencies introduced by drastic viewpoint differences. These inconsistencies disrupt the alignment between positive pairs and ultimately limit the drone-satellite matching performance.
@@
In this paper, we revisit the DVGL task from a manifold learning perspective by modeling the feature space of cross-view images as a composite manifold jointly governed by content and viewpoint factors. As illustrated in Fig. 1 (left), prior methods can be interpreted as learning representations on a single composite manifold $\mathcal{M}$ where both factors are entangled. Such entanglement injects viewpoint-induced conflicts in the learned representations, undermining the robustness of contrastive alignment. To address this limitation, we propose CVD (Content-Viewpoint Disentanglement), a general DVGL framework that explicitly factorizes the feature space into two submanifolds: content $\mathcal{M}_{c}$ and viewpoint $\mathcal{M}_{v}$ (see Fig. 1 (right)). The content encodes view-agnostic geo-structural information, while the viewpoint captures view-specific appearance variations. As shown in Fig. 2, CVD adopts an embed-disentangle-reconstruct paradigm: each image is first embedded into a shared feature space, then projected onto independent content and viewpoint submanifolds, and finally recombined via an image reconstruction task. To facilitate the disentanglement, we impose two dedicated constraints: an intra-view independence constraint and an inter-view reconstruction constraint. The former encourages statistical independence between content and viewpoint by minimizing their mutual information, while the latter preserves the intended semantics of each factor by reconstructing one view using the content from the drone (or satellite) image and the viewpoint from its satellite (or drone) counterpart. In addition, we apply standard contrastive loss (e.g., InfoNCE) to align content representations of matched drone-satellite pairs.
To the best of our knowledge, CVD is the first DVGL framework to explicitly disentangle content and viewpoint. Unlike approaches that require bespoke architectural redesigns, CVD integrates as a plug-and-play module into existing pipelines and lowering inference-time overhead. Extensive experiments on four benchmarks, i.e. University-1652, SUES-200, CVUSA and CVACT, demonstrate consistent gains over multiple baselines, improving robustness to viewpoint and altitude changes and generalization under scene shifts. In summary, our contributions are as follows:
@@
  • We revisit the DVGL task from a manifold-learning perspective and propose CVD, the first framework that explicitly disentangles content and viewpoint to suppress viewpoint-induced conflicting information.
@@
  • We introduce two constraints to facilitate disentanglement: an intra-view independence constraint that facilitates the independence between content and viewpoint, and an inter-view reconstruction constraint that ensures each factor preserves its intended semantics
@@
  • CVD integrates seamlessly into existing pipelines, shortens inference time, and consistently improves cross-view matching performance, enabling efficient deployment in practice.
  • Extensive experiments show that CVD improves the robustness and generalization of DVGL pipelines across diverse scenarios, viewpoints, and altitudes.
@@ 2.1 Visual-based Geo-Localization
Visual-based geo-localization (VGL) has witnessed significant progress with the availability of large-scale geo-tagged datasets and advances in deep learning. Existing methods predominantly follow a Siamese-based framework and can be broadly categorized into three research directions: data augmentation strategies, architectural innovations, and feature representation learning.
@@
Data Augmentation Strategies. Data augmentation has become a widely adopted strategy in VGL tasks [vo2016localizing, rodrigues2021these, zhang2024benchmarking, li2024unleashing]. To address cross-view misalignment, Liu et al.[liu2019lending] incorporate camera orientation as auxiliary input, while SAFA [shi2019spatial] applies a polar transformation to align aerial and ground-level panoramas. CVGlobal [ye2024cross] introduces a panoramic BEV transformation based on the ground-plane assumption and geometric constraints, effectively reducing the gap between street panoramas and satellite imagery. Similarly, Video2BEV [ju2025video2bev] transforms drone videos into BEV representations, facilitating better alignment with satellite imagery. More recently, training-aware data sampling has emerged as a complementary strategy. Sample4Geo [deuser2023sample4geo] introduces two curriculum-driven strategies: one leveraging geographically adjacent samples for easier early-stage alignment, and another mining hard negatives to refine the decision boundary. DenseUAV [dai2023vision] integrates metric learning with mutual supervision, effectively reducing modality discrepancy and improving feature discriminability. Game4Loc [ji2025game4loc] proposes a mutual-exclusion sampling mechanism that enforces strict decorrelation between positive and negative pairs, thereby enhancing contrastive supervision in cross-view matching.
@@
Backbone Innovations. In parallel, a line of work focuses on designing more powerful visual backbones to enhance localization performance [rodrigues2022global, zhu2023modern]. For example, L2LTR [yang2021cross] exploits self-attention to model long-range dependencies, effectively reducing visual ambiguity in cross-view. RK-Net [lin2022joint] introduces a lightweight unit-difference attention module that enables joint learning of dense features and salient keypoints, without requiring additional annotations. SAIG [zhu2023simple] proposes an efficient backbone tailored for VGL by replacing the MLP blocks in standard Transformers with spatially-aware mixing layers and low-dimensional projections, yielding a more compact and structured representation.
@@
Feature Representation Learning. Many studies focus on learning more effective visual representations to enhance cross-view matching [zhu2021vigor, mi2024congeo, zhu2023sues, zhai2017predicting, zeng2022geo, ti2024tirsaa]. A common strategy involves refining alignment mechanisms between views. For instance, Shi et al.[shi2020looking] proposed a dynamic similarity matching network to estimate directional alignment, thereby reducing cross-view discrepancies. FSRA [dai2021transformer] leverages transformer-based heatmaps to perform region-level alignment, while LPN [wang2021each] incorporates contextual cues via a square-ring partitioning strategy to improve part-based representations. SDPL [chen2024sdpl] builds upon LPN by introducing a shifting-fusion mechanism to generate multiple complementary part sets, which are then adaptively aggregated to enhance robustness against spatial shifts and scale variations. Several methods also incorporate inductive priors modeling to enhance feature expressiveness. FRGeo [zhang2024aligning] enhances cross-view alignment by explicitly recombining spatial features to reduce geometric ambiguities. TransGeo [zhu2022transgeo] adopts a non-uniform cropping strategy that discards low-information regions while reallocating resolution to semantically salient areas, enhancing accuracy without increasing computational cost. MCCG [shen2023mccg] enriches feature diversity by jointly modeling spatial and channel-wise attentions.
@@
Although prior methods perform well across various scenarios, many rely on auxiliary components that increase inference-time overhead. In contrast, we propose a plug-and-play training paradigm that explicitly disentangles content and viewpoint from raw feature representations, thereby effectively enhancing cross-view correspondence and reducing inference latency and computational cost.
2.2 Disentangled Representation Learning
Disentangled representation learning (DRL) aims to learn representations that identify and disentangle the underlying factors hidden in observable data [wang2024disentangled]. Owing to the resulting interpretability, controllability, and robustness, it has seen broad adoption in computer vision [li2025fd2, ruan2022adaptive, wang2016multimodal, chen2024disendreamer, cheng2024disentangled], natural language processing [zhao2024disentangled, bao2019generating, cheng2020improving], recommender systems [ma2019learning], and graph learning [wang2020disentangled, zhang2023dyted, wang2022disencite], with gains on many downstream tasks. For example, Zou et al. [zou2020joint] tackle cross-domain person re-identification by jointly disentangling ID-related and ID-unrelated subspaces and restricting adaptation to the former, thereby improving transferability. DisCo [Wang_2024_CVPR] introduces a disentangled-control architecture that separates subject, background, and pose, enabling compositional and generalizable dance video synthesis. Wang et al. [wang2024towards] present a frequency-domain disentanglement framework for UAV object detection that employs two learnable filters to isolate domain-invariant from domain-specific spectra, leading to stronger domain generalization.
Recent studies have begun to incorporate DRL into VGL tasks. GeoDTR [zhang2023cross] introduces a geometry-aware layout extractor to separate geometric cues from raw appearance features, thereby improving cross-view localization. However, it leaves unaddressed the viewpoint-induced conflicts persisting in both appearance and layout representations, thereby hindering cross-view correspondence. In this work, we explicitly disentangle content from viewpoint, and employ independence and reconstruction constraints to suppress such conflicts, yielding a cleaner, view-agnostic content representation.
3 Methods
3.1 Problem Formulation
Considering a set of image pairs $\{(\mathcal{I}_{i}^{\mathrm{d}},\mathcal{I}_{i}^{\mathrm{s}})\}_{i=1}^{N}$, where superscripts $\mathrm{d}$ and $\mathrm{s}$ denote drone and satellite images, respectively, and $N$ is the number of pairs. Each pair depicts the same geographic location. In the DVGL task, given a drone image $\mathcal{I}_{d}^{\mathrm{d}}$ with index $d$, the objective is to retrieve its best-matching satellite image $\mathcal{I}_{b}^{\mathrm{s}}$ from the georeferenced database, where $b\in\{1,\ldots,N\}$.
For effective comparison between cross-view images, we aim to modulate raw representations by explicitly disentangling two factors: content and viewpoint. Concretely, we model the feature space $\mathcal{Z}$ as a representation of composite manifold $\mathcal{M}$ structured by two independent submanifolds, $\mathcal{M}_{c}$ and $\mathcal{M}_{v}$, corresponding to content and viewpoint, respectively. Each feature representation is viewed as a sample from two latent random variables, $C\sim p(C)$ and $V\sim p(V)$, defined over $\mathcal{M}_{c}$ and $\mathcal{M}_{v}$. These are composed via a function $f:\mathcal{M}_{c}\times\mathcal{M}_{v}\rightarrow\mathcal{Z}$ that maps both factors to a point in the feature space. Assuming statistical independence between the factors, i.e., p(C,V)=p(C)p(V)p(C,V)=p(C)\,p(V), the resulting distribution over $\mathcal{Z}$ is given by the push-forward measure $f_{\#}(p(C)\times p(V))$. This assumption is well-aligned with the DVGL task, where identical scenes may be observed under diverse perspectives.
To realize the above formulation, we design manifold encoders that decompose each image $\mathcal{I}^{\mathrm{u}}$ into two distinct representations: a content embedding $f_{c}(\mathcal{I}^{\mathrm{u}})\in\mathcal{M}_{c}$ and a viewpoint embedding $f_{v}(\mathcal{I}^{\mathrm{u}})\in\mathcal{M}_{v}$. Cross-view matching is subsequently performed in the $\mathcal{M}_{c}$ by retrieving the nearest neighbor of a drone image $\mathcal{I}_{q}^{\mathrm{d}}$ via its content embedding:
$b=\underset{i\in\{1,\ldots,N\}}{\arg\min}d(f_{c}(\mathcal{I}_{d}^{\mathrm{d}}),f_{c}(\mathcal{I}_{i}^{\mathrm{s}})),$ (1)
where $d(\cdot,\cdot)$ denotes a distance metric. For notation compactness, we will use superscript uu [^2] We adopt this convention throughout the paper. for cases that apply to both drone ($\mathrm{d}$) and satellite ($\mathrm{s}$) views.
3.2 Proposed Methodology
As illustrated in Fig. 2, CVD adopts a Siamese architecture consisting of two symmetric branches for the drone and satellite views. Each branch comprises three sequential components: manifold embedding, information disentanglement, and cross-reconstruction.
Manifold Embedding. Each input image $\mathcal{I}^{\mathrm{u}}$ ($u\in\{\mathrm{d},\mathrm{s}\}$) is first processed by manifold encoders $E$, yielding a raw $d$-dimensional feature representations $\mathbf{z}^{\mathrm{u}}=E(\mathcal{I}^{\mathrm{u}})\in\mathbb{R}^{d}$. Owing to the nature of DNN encoders, the distribution of $\mathbf{z}^{\mathrm{u}}$ can be viewed as residing on a composite manifold jointly governed by content and viewpoint information. Since our primary focus is on the training paradigm, we adopt the same DNN encoders as those used in the respective baselines to ensure fair and consistent comparisons.
@@ 3.2 Proposed Methodology
Information Disentanglement. Once we obtain the feature representation $\mathbf{z}^{\mathrm{u}}$, we project it into two statistically independent components using two parallel $3\times 3$ convolutional layers with a channel ratio of $\alpha$. One is used to represent the content embedding, denoted as $\mathbf{z}_{c}^{\mathrm{u}}=f_{c}(\mathbf{z}^{\mathrm{u}})\in\mathbb{R}^{\alpha d}$, which provides view-agnostic scene structure. The other is employed for the viewpoint embedding, denoted as $\mathbf{z}_{v}^{\mathrm{u}}=f_{v}(\mathbf{z}^{\mathrm{u}})\in\mathbb{R}^{(1-\alpha)d}$, capturing view-specific attributes. To promote effective disentanglement, we introduce an intra-view independence constraint (Sec. 3.3) that minimizes the mutual information $\mathrm{MI}(Z_{c}^{\mathrm{u}};Z_{v}^{\mathrm{u}})$, where $Z^{\mathrm{u}}$ denotes the random variables of $\mathbf{z}^{\mathrm{u}}$, thereby encouraging statistical independence between content and viewpoint factors.
@@
Cross-Reconstruction. While the independence constraint enhances factor separation, it may inadvertently lead to degenerate solutions or information loss. To mitigate this, we introduce an inter-view reconstruction constraint (Sec. 3.4) that encourages each factor to retain its intended information through cross-view image reconstruction. Specifically, we train two decoders, $D^{\mathrm{d}}$ and $D^{\mathrm{s}}$, to perform bidirectional reconstruction between paired views by swapping content and viewpoint embeddings, i.e., reconstructing each image using its own viewpoint and the content of the other. By enforcing accurate reconstruction from these hybrid embeddings, the model is incentivized to encode factor-specific information in each representation. This cross-view supervision not only prevents information collapse but also reinforces disentanglement. In the following, we describe the two specific constraints in CVD.